# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### 12 Jan 2026 v2

@@ 1 Abstract
Scientific theories of consciousness should be falsifiable and non-trivial. Recent research has given us formal tools to analyze these requirements of falsifiability and non-triviality for theories of consciousness. Surprisingly, many contemporary theories of consciousness fail to pass this bar, including theories based on causal structure but also (as I demonstrate) theories based on function. Herein I show these requirements of falsifiability and non-triviality especially constrain the potential consciousness of contemporary Large Language Models (LLMs) because of their proximity to systems that are equivalent to LLMs in terms of input/output function; yet, for these functionally equivalent systems, there cannot be any falsifiable and non-trivial theory of consciousness that judges them conscious. This forms the basis of a disproof of contemporary LLM consciousness. I then show a positive result, which is that theories of consciousness based on (or requiring) continual learning do satisfy the stringent formal constraints for a theory of consciousness in humans. Intriguingly, this work supports a hypothesis: If continual learning is linked to consciousness in humans, the current limitations of LLMs (which do not continually learn) are intimately tied to their lack of consciousness.
2 A Falsifiability Approach to Consciousness
@@
In the past decade, a subfield of consciousness research has taken a “structural turn” and become focused on mathematics and formalisms [48, 49]. Much of this has stemmed from the influence of Integrated Information Theory (IIT) [2] becoming a popular theory of consciousness, since it is expressed in a mathematically formalized way [11, 47]. In turn, many of the criticisms of IIT have depended on its detailed particulars and mathematical formalisms and whether it is scientifically testable [12, 32]. These debates have led to increased interest in defining things like minimal “toy models” of consciousness in order to differentiate potential theories and test their boundaries [38, 4], and identify what sort of requirements might constrain theories of consciousness.
@@
This structural turn gives rise to a historically novel question: Can a theory of consciousness be so strongly constrained by requirements for falsifiability and non-triviality that its overall nature can be deduced?
@@
Increasingly, such an approach seems a necessity. Right now, the field of consciousness research has hundreds of existing theories [65, 54], and no agreed-upon way to empirically distinguish them. The search for the neural correlates of consciousness [27, 51] has been revealed to have a serious problem wherein even tightly-controlled and well-funded adversarial collaborations fail [26], such as the recent comparison of IIT and Global Neuronal Workspace Theory [57, 9], which led to accusations of pseudoscience [32, 67]. If consciousness research is to make serious progress in our lifetime, it must winnow the field via constraints (see, e.g., [30]).
Compared to other approaches, working backwards from “What kinds of theories of consciousness are falsifiable and non-trivial?” is a far less explored option for deriving a class of viable theories. In this, scientists would act like artists drawing the negative space around consciousness to see its outline clearly, with minimal metaphysical assumptions and no adherence to any particular theory.
@@
Another reason to pursue this falsification-first approach is that we want questions about consciousness that the contemporary state of consciousness research cannot answer yet. Most prominently, the question whether or not contemporary LLMs (like ChatGPT or Claude or Gemini) are conscious has become suddenly critical. There are major risks associated with getting this question wrong. The more scientifically conservative perspective is that contemporary LLMs are not conscious and yet may (depending on prompt) seem conscious and so be taken as if conscious. This has a myriad of risks, which include increasing the risk of AI psychosis, overestimation of their capabilities and import, inappropriate practices or regulation, and misleading scientific beliefs about human consciousness [66, 58]. On the other hand, if contemporary LLMs were conscious they could be considered moral patients [64] and be due considerations like conversational “exit rights” [7] as well as other potential rights against mistreatment.
@@
It seems unlikely that empirical data from LLMs alone can answer the question of their consciousness. LLMs are highly prone to confabulation [41]. And while there is a small subfield studying LLM “introspection” [14], the best evidence for this introspection is in the form of statistically uncommon anomaly detection of injected concepts, for which there could be shallow mechanistic explanations [8]. There is not even agreement over whether (low-resolution) memory stored from a forward pass of previous internal processing actually counts as anything like human “introspection,” nor if it should be associated with consciousness.
@@
Therefore, most have assumed that answering the question of LLM consciousness requires a scientific understanding of consciousness beyond the current day (e.g., scientific consensus around a particular theory of consciousness). However, I show that this is not necessary. Instead, it is possible to prove that there is no theory of consciousness that is non-trivial and falsifiable that could exist for LLMs. This represents much faster progress than trying to assign assumption-dependent probabilities of consciousness, or trying to apply some set of the many (often contradictory) theories that do currently exist [20].
@@
This road to a disproof of LLM consciousness begins with one particularly influential criticism of IIT: The “Unfolding Argument” by Doerig et al. [31] has led to an evocative debate in the literature on precisely how to falsify theories of consciousness. Specifically, the argument is based on the universal approximation theorem [39], and argues that any given Recurrent Neural Network (RNN) can be “unfolded” into some functionally equivalent (in terms of input/output) Feedforward Neural Network (FNN). This creates a pathological scientific situation for theories like IIT, since any given RNN might have a certain degree of integrated information, yet its input/output “twin” FNN would have zero integrated information.
@@
However, the Unfolding Argument alone only brings to light that internal “causal structure” theories like IIT face significant challenges. In later work by Johannes Kleiner and myself [45], we reached two further conclusions. First, that unfolding being possible was a consequence of the independence between what a theory uses to make predictions about consciousness (like the integrated information of IIT) and how a scientist infers the actual states of conscious experiences during experiments (via things like report or behavior or more broadly, outputs). We argued that this made the unfolding argument apply to a broader swath of theories than initially thought, and that the best conceptualization of the issue is a framework wherein alternative systems can be “substituted” in during empirical testing in order to reveal pathologies around falsifiability. Additionally, we pointed out that there is also a risk on the other side (the opposite problem) of having too much dependency between a theory’s predictions and a scientist’s inferences, which would create theories of consciousness that are trivial and unfalsifiable. In other words, there are two “horns” to the problem of falsifiability for theories of consciousness.
The dual-horned “Kleiner-Hoel dilemma” established in [45] goes like this: A successful theory of consciousness must not be a priori falsified by the class of arguments like the Unfolding Argument (broadly, any such “universal substitution,” like unfolding, that creates a pathological circumstance of changing all, or at least the vast majority, of predictions, while holding inferences constant). However, the same theory must also not be unfalsifiable in principle. This creates only a small space through which to pass; theories of consciousness are “caught between Scylla and Charybdis.”
@@
On one side, theories like IIT [68] or recurrent processing theories [56] are a priori falsified by substitutions like unfolded neural networks (since pretty much all the predictions of a theory would change under such a substitution, while the experimenter’s inferences based on reports or behavior would stay exactly the same, creating universal mismatches between predictions and inferences and thus resulting in falsification). Meanwhile, theories like behaviorism or global functionalism (both being based solely on input/output mattering for consciousness) are unfalsifiable, because predictions of the theories and inferences about consciousness from experiments can never differ (e.g., if your theory of consciousness is that consciousness is identical to behaving as if conscious, how could it ever be falsified?).
This was a strong challenge to the consciousness literature. The primary reply against it has been simply that the Kleiner-Hoel dilemma is too strong and therefore cannot possibly be true, at least without ruling out consciousness research in general [72]. This is a debatable reason to reject a conclusion, but more importantly, it falls by the wayside once an appropriate theory is introduced (even a sketch of one) that can navigate the Kleiner-Hoel dilemma. Critically, if the Kleiner-Hoel dilemma is navigated successfully in humans, but provably cannot be navigated in LLMs (due to their close resemblance to systems that we know cannot be targets for non-trivial and falsifiable theories), this would form a strong argument against LLM consciousness. In other words, there may be properties of human brains that allow them to “pass through” the Kleiner-Hoel dilemma unscathed, but LLMs, lacking those properties, cannot.
That is what this paper shows. It is structured in three parts.
@@
First, I overview a formal framework for testing theories of consciousness [45], and demonstrate how the Kleiner-Hoel dilemma naturally arises from it, including a new proof based on Kolmogorov’s notion of a “shortest program” for how even functionalist theories face challenges from substitutions (or fall into triviality).
@@
Second, I show that the Kleiner-Hoel dilemma has extremely strong argumentative force in LLMs, because LLMs are much closer than human brains in “substitution space” to input/output equivalent substitutions like lookup tables that non-trivial theories of consciousness can’t possibly judge as conscious. That is, there is no theory of consciousness available for contemporary LLMs that would not fall on one of the horns of the Kleiner-Hoel dilemma. Therefore, there is no scientific theory of consciousness which could judge contemporary LLMs as conscious. Therefore, they are not conscious.
@@
The strength of this argument is that it is entirely agnostic to any assumptions about particular theories of consciousness being true, nor does it make assumptions about the “specialness” of biology, or the possibility of AI consciousness in general, and has few philosophical commitments (other than that theories of consciousness should be falsifiable and non-trivial).
@@
Third, this paper shows that the Kleiner-Hoel dilemma has a potential solution in humans by grounding a theory of consciousness not in causal structure like IIT, nor functional or dynamical features like Global Workspace Theory, but in continual learning. This fits well with the recent observation by O’Reilly-Shah et al. that a static FNN cannot act as a universal substitution for a plastic RNN [59]. However, those recent proofs only show that plastic RNNs are protected from the specifics of the Unfolding Argument—they do not fully distinguish whether the recurrence or plasticity is the cause, nor ask if this protection extends to other universal substitutions. Here I show that grounding a theory of consciousness in the broader context of continual learning explicitly avoids both horns of the Kleiner-Hoel dilemma.
@@
This result implies that scientific theories of consciousness are, until other cases that avoid the Kleiner-Hoel dilemma are identified, strongly constrained to not merely being confined to systems that can learn (like plastic RNNs), but more generally that we can constrain theories of consciousness to, at minimum, involve continual learning itself. Such a result significantly winnows the field of consciousness research and illuminates the path toward less popular, but ultimately more scientific, theories of consciousness. It also implies that the vast majority of the search for the neural correlates of consciousness may be incommensurate, because testing static input/output behavior alone is not how one empirically tests for consciousness.
@@ 3 A Formal Framework for Falsifying Theories of Consciousness
The line of reasoning here will make use of a previously developed framework for falsification of theories of consciousness [45]. This is a formalized mathematical framework for testing theories of consciousness, based on treating theories as mapping a space of possible experiences onto physical predictions. Its advantage is that it is abstract, general, and (relatively) simple.
Here, I leave much of the background logic and reasoning to previous work and deal only with the high-level details that are relevant (see [45] for those details). For our purposes what matters is that testing a theory of consciousness is based on comparing two functions: predictions and inferences.
@@
Theories are defined by their unique methods of making predictions. However, predictions alone cannot falsify a theory. They need to be compared to empirical inferences about consciousness, either explicit or implicit.
Usually, these inferences are based on verbal report, behavior, or, more recently, even “no-report paradigms” [70]. But even “no-report paradigm” cases (based in things like eye movement) rely on correlating with report first. In a psychology lab, button presses to indicate what was flashed in a masking stimulus are inference data; in another context, like in assessing LLM consciousness, the outputs of an artificial neural network could count as inference data. Inferences are, essentially, empirical reasons to believe a system might be conscious irrespective of any given theory or class of theories. Perhaps the clearest inference would be an explicit claim by the system to consciousness itself, but individual examples of inferences while testing a theory of consciousness might be, e.g., correctly reporting sensory inputs, responding to pain, etc.
Overall, inferences can be represented by another function, which maps between the data generated by observations about the behavior of a system (like its input/output) and the space of experiences $E$:
@@
It is important that a falsification framework not be overfitted or over-sensitive. For the purposes herein, I do not require that falsification triggers from just any possible mismatch between pred and inf. For instance, imagine a theory of consciousness that performs very well in humans and we are starting to believe scientifically. Then, we apply it to some systems for which we cannot access consciousness as well (say, plants). This gives us a (to many) surprising result: that plants are conscious to some degree. However, it may be that no experimental inference from plants can fully support the weight of the claim of the theory, and therefore, there are technically some mismatches. The point here is not to say this should happen with a theory, or is totally non-problematic, but just that, at this time, the falsification framework introduced herein can be optionally permissive about some occasional mismatches.
On the other hand, if, e.g., a theory of consciousness predicts that extremely simple systems have degrees of consciousness greater than that of a human [1, 69], such a radical mismatch could count as actual falsification—but only in light of the inference that a system’s behavioral or input/output simplicity is evidence against such a high degree of consciousness. Ideally, such inferences should always be made explicit when reasoning about consciousness through the lens of falsification. However, it is important to note that keeping assumptions about inferences as neutral as possible is important, or else testing a theory of consciousness would lose its desired empirical “universality” [42].
@@
Importantly, it is obvious that if inf is held constant, but pred is varied significantly, then this will create consistent divergence (mismatches) across the entirety (or vast majority) of a theory’s predictions. This is pathological behavior and should necessarily falsify a given theory, as now every prediction is wrong (or the vast majority are wrong) for a system with identical (or, again, extremely similar) inferences. Specifically, in [45] we took the approach that a “universal substitution” (of one system for another) in an experimental setup (within sensible allowances), but in which inferences do not change—but all predictions do change—would completely falsify a particular theory. We called such a possibility a “universal substitution” and confined the analyses to such cases in order to avoid more debatable instances. E.g., it is unclear to what degree “partial substitutions” triggering mismatches should count as falsification, since it may make sense to sometimes err on the side of a well-confirmed theory in edge cases (but again, this question can be set aside here).
However, what is clearly problematic scientifically are universal substitutions wherein all predictions of a theory are now different following the substitution, but inferences have remained just the same (or, slightly more permissively, the overwhelming majority of inferences remain the same and the overwhelming majority of predictions change). The power of the original Unfolding Argument is precisely that it is a clever way of affixing the input/output of a neural network but implicitly varying all (or, again, the overwhelming majority) of the predictions of a theory of consciousness for the exact same input/output behavior.
As we showed, substitutions are guaranteed to exist as long as predictions and inferences are totally independent. Another way of saying this: Just from knowing the structure of Integrated Information Theory, and that the calculation of $\Phi$ is entirely screened off from the input/output of a system, there can be systems wherein $\Phi$ differs but the input/output does not. However, the reason why this appears to be such a problem for IIT is simply that IIT is formally well-developed. Other theories of consciousness will eventually become equally formally developed, and have to specify similarly precise predictions to $\Phi$, and then they will fall victim to the same problems.
@@
  • Causal-structure substitutions accomplish a particular input/output function via a different system with different internal workings. This includes lookup tables that implement the same input/output function of a system directly [63], or static FNNs that implement the same input/output (or approximate it arbitrarily well) [31]. This might include potentially more exotic options, like substitutions based on the AIXI model of universal intelligence [40] or even cases of a brain in quantum superposition (a “Schrödinger’s Zombie”) [18].
  • Functional substitutions accomplish any particular input/output function via a different algorithm or program with different workings. For instance, there are various ways to sort (bubble sort, merge, etc.) that can all transition an unsorted list to a sorted one.
It had been hoped that issues of falsification like the Unfolding Argument would only ever impact “causal structure” theories like IIT [31]. Alternatively, some have speculated that if this form of argument impacts functionalist theories, it would form a reductio and actually argue against the Kleiner-Hoel dilemma being a problem [72]. However, that its widespread negative result is too strong would only be true if the Kleiner-Hoel dilemma provably cannot be navigated by any theory whatsoever. If a theory (or a class of theories) does navigate it, we cannot say that the requirement is too strong: rather, it is precisely strong enough to substantially constrain consciousness research and point it in the right direction.
For these purposes, in what follows I show precisely how the Kleiner-Hoel dilemma impacts all functionalist theories of consciousness too. Instead of substituting systems, we can substitute programs. Specifically, the shortest program.
@@
Step 2. Let $N$ be any program computing $f$ that is the target of the predictions of some theory of consciousness. $N$ and $K(f)$ share input/output (defined by $f$).
@@
Many theories could be shown to have problems with falsification in this light. Consider just one: Higher Order Thought (HOT) theory, a popular theory of consciousness [19]. Any exact ranking or rating of the different representations should differ in its predictions about $N$ vs. $K(f)$. But if somehow not, then $N$ is irrelevant and only the input/output and its subsequently determined $K(f)$ is needed for the theory’s predictions, rendering it trivial. The same goes for almost all theories of consciousness that are more focused on functionalist aspects, like “representation” or “computation.”
The overall structure of this proof also demonstrates how the Kleiner-Hoel dilemma has two horns. Here, if substituting in the shortest program that computes the function $f$ makes a theory’s predictions change, then it creates mismatches: A kind of “a priori falsification.” However, on the other hand, if the predictions of a theory of consciousness do not change, then they must be strictly dependent on the experimental inferences observed during testing a theory of consciousness, rendering that theory trivial (here, this can be clearly seen by any details of the original program $N$ no longer mattering at all if a theory’s predictions are always the same for $K(f)$ too). This would be a case of “strict dependency” wherein a predictions can be entirely derived by the inferences (in the terminology of Theorem 3.1, inferences are entirely determined by $f$).
@@
Here and throughout, such strictly-dependent-on-inferences theories of consciousness are referred to as “trivial” to clearly distinguish that the source of the unfalsifiability is coming from strict dependency holding true. E.g., a widely-held (but trivial) theory of consciousness is a kind of behaviorism, such that if an entity acts conscious, then it is conscious. But such a theory cannot be falsified, since its predictions could never be any different than experimental inferences (as they are simply a direct restatement of them) and so such a theory contains no scientific information, and so it is trivial. This applies to more than just behaviorism about consciousness. While the term “functionalism” has many different meanings, for the same reason that behaviorist theories of consciousness are trivial, so is any theory of consciousness based solely on input/output. Arguably, this does significantly affect plenty of existing theories, which turn out to be simply behaviorist theories with multiple steps.
E.g., when discussing Global Neuronal Workspace Theory (GNWT), Daniel Dennett once wrote that:
@@
If Dennett’s ideas were true, this would make GNWT a trivial theory of consciousness, since it would be unfalsifiable, because the predictions and inferences strictly depend on the same source. Consciousness just is the information that is globally accessible for report and behavior, and reports and behaviors also just are different expressions of that same information. So there is no situation wherein they could meaningfully diverge. And therefore, the theory is unfalsifiable and trivial (one could see how it gives us minimal scientific information even if it were true—trivial theories are fundamentally unsatisfying).

### 19 Jan 2026 v3

@@ 1 Abstract
Scientific theories of consciousness should be falsifiable and non-trivial. Recent research has given us formal tools to analyze these requirements of falsifiability and non-triviality for theories of consciousness. Surprisingly, many contemporary theories of consciousness fail to pass this bar, including theories based on causal structure but also (as I demonstrate) theories based on function. Herein, I show these requirements of falsifiability and non-triviality especially constrain the potential consciousness of contemporary Large Language Models (LLMs) because of their proximity to systems that are equivalent to LLMs in terms of input/output function; yet, for these functionally equivalent systems, there cannot be any falsifiable and non-trivial theory of consciousness that judges them conscious. This forms the basis of a disproof of contemporary LLM consciousness. I then show a positive result, which is that theories of consciousness based on (or requiring) continual learning do satisfy the stringent formal constraints for a theory of consciousness in humans. Intriguingly, this work supports a hypothesis: If continual learning is linked to consciousness in humans, the current limitations of LLMs (which do not continually learn) are intimately tied to their lack of consciousness.
2 Constraining theories of consciousness to testable theories
@@
In the past decade, a subfield of consciousness research has taken a “structural turn” and become focused on mathematics and formalisms [60, 62]. Much of this is the influence of Integrated Information Theory (IIT) [2], since it is expressed in a mathematically formalized way [10, 59]. In turn, many of the criticisms of IIT have depended on its detailed particulars and whether it is scientifically testable [11, 39]. These debates have led to increased interest in defining minimal “toy models” of consciousness [47, 4] and identifying what requirements might constrain theories of consciousness in terms of being testable (such as being falsifiable).
@@
This structural turn gives rise to a novel approach: Can a theory of consciousness be so strongly constrained by requirements for testability that its overall nature can be deduced?
@@
Increasingly, such an approach seems a necessity. Currently, the field of consciousness research has hundreds of existing theories [83, 67], and no agreed-upon way to empirically distinguish them. The search for the neural correlates of consciousness [31, 64] has revealed that even tightly-controlled and well-funded adversarial collaborations fail [30], such as the recent head-to-head comparison of IIT and Global Neuronal Workspace Theory [71, 8], which led to accusations of pseudoscience [39, 86]. If consciousness research is to make serious progress, it must winnow the wide field via constraints on properties of theories [57, 36]. In this, scientists would act like artists drawing the negative space around consciousness to see its outline clearly.
@@
A further reason to pursue a testability-first approach is that we want questions about consciousness that the contemporary state of consciousness research cannot answer yet. Most prominently, the question whether or not contemporary LLMs (like ChatGPT or Claude or Gemini) are conscious has become suddenly critical. There are major risks associated with getting this question wrong. Assigning consciousness where there is none has a myriad of risks, which include increasing the risk of AI psychosis, overestimation of LLM capabilities, inappropriate practices or regulation, and misleading scientific beliefs about human consciousness [85, 74]. On the other hand, if contemporary LLMs were conscious they could be considered moral patients [82] and be due considerations like conversational “exit rights” [7] as well as other protections against mistreatment.
@@
It is unlikely that empirical data from LLMs alone can answer the question of their consciousness, given how prone they are to confabulation and prompt-sensitivity [50]. And while there is a small subfield studying LLM “introspection” [14], the best evidence for this introspection is in the form of statistically uncommon anomaly detection of injected concepts, for which there could be shallow mechanistic explanations [70].
@@
Therefore, most have assumed that answering the question of LLM consciousness would require scientific consensus around a particular theory of consciousness. However, I show that this is not necessary. Instead, it is possible to prove that there is no theory of consciousness that is non-trivial and falsifiable that could exist for (at minimum) baseline LLMs. This represents much faster progress than trying to assign assumption-dependent probabilities of consciousness, or trying to apply some set of the many (often contradictory) theories that do currently exist [20].
@@
This road to a disproof of LLM consciousness begins with one particularly influential criticism of IIT: The “Unfolding Argument” by Doerig et al. [37] has led to debate in the literature on how to falsify theories of consciousness. Specifically, the argument is based on the universal approximation theorem [48], and argues that any given Recurrent Neural Network (RNN) can be “unfolded” into some functionally equivalent Feedforward Neural Network (FNN). This creates a pathological scientific situation for theories like IIT, since any given RNN might have a certain degree of integrated information, yet its “twin” FNN would have zero integrated information. This is despite the two networks sharing the same input/output (I/O).
@@
Later work by Johannes Kleiner and myself [57] showed that unfolding was a specific case of the independence between what a theory uses to make predictions about consciousness (like the integrated information of IIT) and how a scientist infers the actual states of conscious experiences during experiments (via things like report or behavior or, more broadly, any I/O). We proposed a framework wherein alternative systems (of any kind) can be “substituted” in during empirical testing while holding I/O fixed, which reveals pathologies around falsifiability (thus, the ”Substitution Argument”).
Additionally, we pointed out that, instead of independence, there is a risk of having too much dependency between a theory’s predictions and a scientist’s inferences. This would create theories of consciousness that are trivial. In other words, there are two “horns” for theories of consciousness: the Substitution Argument on one side, and triviality on the other. The dual-horned “Kleiner-Hoel dilemma” established in [57] creates only a small space through which to pass; theories of consciousness are “caught between Scylla and Charybdis.”
@@
According to the Substitution Argument, theories like IIT [87] or recurrent processing theories [69] are a priori falsified by specific substitutions like unfolded neural networks [37]. Meanwhile, theories like behaviorism or functionalism based solely on I/O are trivial, because predictions of any such theory can never differ from inferences about consciousness from experiments (which are also based on I/O). E.g., if a theory of consciousness is that consciousness is identical to behaving as if conscious, how could it ever be falsified? It gives only trivial scientific information.
This paper shows that the Kleiner-Hoel dilemma is particularly troublesome for LLMs, and gives a disproof of their consciousness.
@@
First, I overview a formal framework for testing theories of consciousness [57], and demonstrate how the Kleiner-Hoel dilemma naturally arises from it, including a new proof based on Kolmogorov’s notion of a “shortest program” for how even functionalist theories face challenges from substitutions (or, depending on how they’re defined, fall into triviality).
@@
Second, I show that the Kleiner-Hoel dilemma has especially strong force in LLMs, because LLMs are much closer than human brains in “substitution space” to I/O equivalent substitutions like lookup tables that non-trivial theories of consciousness can’t possibly judge as conscious. That is, there is no theory of consciousness available for contemporary LLMs that would not fall on one of the horns of the Kleiner-Hoel dilemma. Therefore, there is no scientific theory of consciousness which could judge contemporary LLMs as conscious. Therefore, they are not conscious.
@@
The strength of this argument is that it is entirely agnostic to particular theories of consciousness being true, nor assumes the “specialness” of biology, nor outright dismisses the possibility of AI consciousness in general, and has few philosophical commitments (other than that theories of consciousness should be falsifiable and non-trivial).
@@
Third, this paper sketches a solution to the Kleiner-Hoel dilemma in humans by grounding a theory of consciousness not in causal structure like IIT, nor functional or dynamical features like Global Workspace Theory, but in continual learning. This fits well with the recent observation by O’Reilly-Shah et al. that a static FNN cannot act as a universal substitution for a plastic RNN [75]. However, those recent proofs only show that plastic RNNs are protected from the specifics of the Unfolding Argument—they do not fully distinguish whether the recurrence or plasticity is the cause, nor ask if this protection extends to other universal substitutions. Here I show that grounding a theory of consciousness in the broader context of continual learning explicitly avoids both horns of the Kleiner-Hoel dilemma.
@@
Assuming that a robust scientific theory of consciousness exists, until other cases that avoid the Kleiner-Hoel dilemma are identified, they are strongly constrained to not merely being confined to systems that can learn (like plastic RNNs), but more generally to involve continual learning itself. Such a result significantly winnows the field of consciousness research and illuminates the path toward less popular, but ultimately more scientific, theories of consciousness. It also implies that the vast majority of the search for the neural correlates of consciousness may be incommensurate, because testing static I/O behavior alone is not how one empirically tests for consciousness.
@@ 3 A Formal Framework for Falsifying Theories of Consciousness
The line of reasoning here will make use of a previously developed framework for falsification of theories of consciousness [57]. It is based on treating theories as mapping a space of possible experiences onto predictions about the system’s workings. Its advantage is that it is abstract, general, and (relatively) simple and broadly follows common philosophical definitions of falsification [79, 68].
I leave much of the background logic and reasoning to previous work (see [57] for those details). For our purposes what matters is that testing a theory of consciousness is based on comparing two functions: predictions and inferences.
@@
Theories are defined by their unique methods of making predictions. However, predictions alone cannot falsify a theory. They need to be compared to empirical inferences about consciousness.
Usually, these inferences are based on verbal report or behavior. In a psychology lab, button presses to indicate what was flashed in a masking stimulus could be inference data; in another context, like in assessing LLM consciousness, the outputs of an artificial neural network could be inference data. Inferences are, essentially, empirical reasons to believe a system might be conscious irrespective of any particular theory or class of theories. Perhaps the clearest inference would be an explicit claim by the system to consciousness itself, but others might be, e.g., correctly reporting sensory inputs, responding to pain, etc.
Inferences can be represented by a second function, which maps between the data generated by observations about the behavior of a system (like its I/O) and the space of experiences $E$:
@@
It is important that a falsification framework not be overfitted or over-sensitive. For the purposes herein, I do not require that falsification triggers from just any possible mismatch between pred and inf. Therefore, here I remain agnostic to whether the falsification framework introduced herein can be optionally permissive about some occasional mismatches.
However, if a theory of consciousness predicts that extremely simple systems have degrees of consciousness greater than that of a human [1, 88], such a radical mismatch could count as actual falsification—but only in light of the inference that a system’s behavioral or I/O simplicity is evidence against such a high degree of consciousness. Ideally, such inferences should always be made explicit when reasoning about consciousness through the lens of falsification. It is important to note that keeping assumptions about inferences as neutral as possible is important, or else testing a theory of consciousness would lose its desired empirical “universality” [52] and become less robust.
@@
Importantly, if inf is held constant, but pred is varied significantly, then this will create consistent divergence (mismatches) across the entirety (or vast majority) of a theory’s predictions. This is pathological behavior and should necessarily falsify a given theory, as now every prediction is wrong (or the vast majority are wrong) for a system with identical (or, again, extremely similar) inferences. Specifically, in [57] we took the approach that a “universal substitution” (of one system for another) in an experimental setup (within sensible allowances), but in which inferences do not change—but all predictions do change—would completely falsify a particular theory. We called such a possibility a “universal substitution” and confined the analyses to such cases in order to avoid more debatable instances. E.g., it is unclear to what degree “partial substitutions” triggering mismatches should count as falsification (but again, this question can be set aside here). However, what is clear is that universal substitutions stem from predictions and inferences being totally independent.
@@
  • Causal-structure substitutions accomplish a particular I/O function via a different system with different internal workings. This includes lookup tables that implement the same I/O function of a system directly [81], or static FNNs that implement the same I/O (or approximate it arbitrarily well) [37]. This might include potentially more exotic options, like substitutions based on the AIXI model of universal intelligence [49] or even cases of a brain in quantum superposition (a “Schrödinger’s Zombie”) [18].
  • Algorithmic substitutions accomplish any particular I/O function via a different algorithm or program with different workings. For instance, there are various ways to sort (bubble sort, merge, etc.) that can all transition an unsorted list to a sorted one.
It had been hoped that issues of falsification like the Unfolding Argument would only ever impact “causal structure” theories like IIT [37]. However, once generalized as the Substitution Argument, there did not seem any particular protection for functionalist or computational theories [57]. This is likely because the definition of what is functionalist and what is causal structure is difficult to define (e.g., functionalism may require a “causal topology” [22]). In what follows I show precisely how the Kleiner-Hoel dilemma impacts functionalist and computational theories of consciousness too. Instead of substituting systems, we can substitute programs. Specifically, the shortest program.
@@
Step 2. Let $N$ be any program computing $f$ that is the target of the predictions of some theory of consciousness. $N$ and $K(f)$ share I/O (defined by $f$).
@@
Many theories could be shown to have problems with falsification in this light. Consider just one: Higher Order Thought (HOT) theory, a popular theory of consciousness [19]. Any exact ranking or rating of the different representations should differ in its predictions about $N$ vs. $K(f)$. But if somehow not, then $N$ is irrelevant and only the I/O and its subsequently determined $K(f)$ is needed for the theory’s predictions, rendering it trivial. The same goes for almost all theories of consciousness that are more focused on functionalist aspects, like “representation” or “computation.” As a concrete example, it has been argued that “Perceiver” architecture of a neural network satisfies the properties of Global Workspace Theory [51, 33]. This architecture could be replaced by a lookup table, a single-hidden-layer FNN, or $K(f)$, all while holding I/O constant, and all involving substantial variations in predictions (such as a lack of consciousness).
The overall structure of this proof also demonstrates how the Kleiner-Hoel dilemma has two horns. The first horn is the Substitution Argument: If substituting in the shortest program that computes the function $f$ makes a theory’s predictions change, then it creates mismatches, enacting a kind of “a priori falsification.” However, on the other hand, if the predictions of a theory of consciousness do not change, then they must be strictly dependent on the experimental inferences observed during testing a theory of consciousness, rendering that theory trivial (here, this can be seen by any details of the original program $N$ no longer mattering at all if a theory’s predictions are always the same for $K(f)$ too). This would be a case of “strict dependency” wherein predictions either mimic, stem from, or can be entirely derived by the inferences (in the terminology of Theorem 3.1, inferences would be entirely determined by $f$).
@@
Here and throughout, such strictly-dependent-on-inferences theories of consciousness are referred to as “trivial.” E.g., a widely-held (but trivial) theory of consciousness is behaviorism, such that if an entity acts conscious, then it is conscious. But such a theory cannot be falsified, since its predictions could never be any different than experimental inferences (as they are simply a restatement of them) and so such a theory contains no scientific information, and so it is trivial. While the term “functionalism” has many different meanings, for the same reason that behaviorist theories of consciousness are trivial, so is any theory of consciousness based solely on I/O. Arguably, this does significantly affect plenty of existing theories, which turn out to be simply behaviorist theories with multiple steps. E.g., when discussing Global Neuronal Workspace Theory (GNWT), Daniel Dennett once wrote that:
@@
If Dennett’s ideas were true, this would make GNWT a trivial theory of consciousness, since it would be unfalsifiable, because the predictions and inferences strictly depend on the same source. Consciousness just is the information that is globally accessible for report and behavior, and reports and behaviors also just are different expressions of that same information. So there is no situation wherein they could diverge. And therefore, the theory is unfalsifiable and trivial (one could see how it gives us minimal scientific information even if it were true—trivial theories are fundamentally unsatisfying).